# Learning to Disentangle Latent Reasoning Rules with Language VAEs: A Systematic Study

### v1

@@ Abstract
Incorporating explicit reasoning rules within the latent space of language models (LMs) offers a promising pathway to enhance generalisation, interpretability, and controllability. While current Transformer-based language models have shown strong performance on Natural Language Inference (NLI) tasks, they often rely on memorisation rather than rule-based inference. This work investigates how reasoning rules can be explicitly embedded and memorised within the LMs through Language Variational Autoencoders (VAEs). We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. Performance bottleneck: In mathematical reasoning tasks using Qwen2.5 (0.5B), increasing sample count doesn’t improve performance beyond a point. Moreover, ffn layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters.
1 Introduction
@@
Encoding reasoning patterns or explicit rules within latent representations holds significant promise for enhancing the generalisation, interpretability, and controllability of neural models, which is particularly salient in the neuro-symbolic AI domain Bonnet and Macfarlane (2024); Yu et al. (2025).
@@
Over recent years, Transformer-based language models (LMs) have achieved notable success across a variety of Natural Language Inference (NLI) tasks Yang et al. (2024); Qwen et al. (2025). Nonetheless, a growing body of research has demonstrated that these models often rely on memorisation rather than rule-based inference mechanisms, thereby constraining their capacity for generalisation and interpretability Yan et al. (2025).
Therefore, this study focuses on the question: How can reasoning rules (i.e., memorised input-output patterns [^1] Please notice that exploring the composition and generalisation over rules in the NLI task is not the target of this work. The motivation of this work is to explicit encode the memorised reasoning input-output patterns in the latent space.) be explicitly encoded within the latent space of language models?
@@
Variational Autoencoder (VAE) Kingma and Welling (2013) provides a compelling framework for this direction. The integration of prior distribution serves as an inductive bias, enabling the model to leverage existing knowledge and providing a principled way to incorporate domain expertise Papamarkou et al. (2024). In this work, we focus on lightweight LMs (<1B parameters) as VAE decoders, allowing accessible evaluation of training dynamics, memory capacity, and model updates of the Transformer architecture Zhong et al. (2025); Morris et al. (2025). More importantly, because our primary motivation is to study latent rule encoding and separation, we fully re-train the decoder. As a result, the geometry and separation of the latent space are not constrained by the model’s size or architectures, enabling a more lightweight and controlled analysis of representation learning.
@@
Accordingly, we propose a complete pipeline for learning NLI reasoning rules within Transformer-based language VAEs Li et al. (2020); Zhang et al. (2024c): First, we focus on three distinct rule-based reasoning tasks, each characterised by unique syntax, inference patterns and levels of granularity: Mathematical Reason Meadows et al. (2023), Syllogistic Reason Valentino et al. (2025), and Explanatory Reason Zhang et al. (2024b).
@@
Second, following the Neural Tangent Kernel (NTK) theory Jacot et al. (2018), we claim that by supervising the inference rule information, the inference rule can be encoded in the parametric space (i.e., weight matrix) of encoder, which subsequently leads to the rule separation in the latent sentence space. This formalisation allows us to explicit encode the reasoning rule in the latent space.
@@
Third, we propose an end-to-end VAE architecture where the latent space can encode both coarse-grained reasoning rules and fine-grained reasoning patterns that arise from semantic differences in the inputs, and evaluate three ways to inject the latent space into the LM decoder by intervening the attention network.
@@
We conduct extensive experiments to assess the effectiveness of rule encoding and reasoning generation capabilities. The experiment illustrates the following results and findings:
@@
Disentangled Reasoning: Under explicit signal supervision, reasoning rules—viewed as functional mappings—can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space, suggesting the potential for employing rule-based NTK theory to better understand the training dynamics and internal geometry of gradient-based neural NLI model.
@@
Prior Knowledge Injection: Different injection setups result in varying levels of reasoning performance. The optimal configuration is achieved by injecting the latent space into the query matrix of the attention network. Intuitively, injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models.
@@
Performance Bottleneck: A case study on the mathematical reasoning task reveals a performance bottleneck in decoder-only LM, Qwen2.5 (0.5B), where increasing the number of samples per operation fails to yield performance improvements beyond a certain threshold. Additionally, we observe that ffn network is more effective than attention in maintaining the separation of reasoning rules within the learned parametric space.
@@
To the best of our knowledge, this is the first study to explore the explicit encoding of NLI rules within the latent space of language VAEs.
2 Related Work
@@
In this section, we review the related work around two topics: rule-based representation learning and language VAEs, to highlight the current research limitation and elucidate the motivation underlying our work.
@@
Encoder-only models have been employed for tasks such as mathematical operations, where the encoder learns structured transformations, as demonstrated by Valentino et al. (2024). In addition, VAE-based approaches have shown promise in tasks requiring structured reasoning, such as visual program synthesis, where the goal is to generate programs that fulfill a specified task Sun et al. (2018); Bonnet and Macfarlane (2024); van Krieken et al. (2025). Similarly, grammar-based approaches using VAEs have been applied to infer ordinary differential equation (ODE) formulas from data Yu et al. (2025). Moreover, Decoder-only models, particularly large language models (LLMs), leverage in-context learning by using demonstrations to infer and apply underlying reasoning patterns Liu et al. (2023b); Bhattamishra et al. (2024). Despite these advancements, relatively few studies have investigated rule-based learning in the context of NLI. To address this gap, we present a foundational exploration of rule-based learning applied to the NLI task.
@@
Language VAEs have been widely applied in NLP tasks, such as style transfer tasks: modifying sentences with regard to markers of sentiment, formality, affirmation/negation Shen et al. (2020); John et al. (2019); Bao et al. (2019b); Hu and Li (2021); Vasilakes et al. (2022); Gu et al. (2022); Liu et al. (2023a); Gu et al. (2023), story generation Fang et al. (2021), dialogue generation Zhao et al. (2017), text paraphrasing Bao et al. (2019a), and textual, syntactic, semantic representation learning domain, such as syntax disentanglement Mercatali and Freitas (2021), semantic-syntax separation Zhang et al. (2024c), semantic disentanglement Silva De Carvalho et al. (2023); Zhang et al. (2024a), etc. Comparatively, we focused on the Natural Language Inference (NLI) task. To our knowledge, the application of VAEs for NLI is underexplored.
@@
3 Rule-based Learning for NLI
3.1 Natural Language Inference Rules
In this study, we investigate three distinct types of reasoning tasks: mathematical derivation Meadows et al. (2024), syllogistic reasoning Valentino et al. (2025), and explanatory reasoning Zhang et al. (2024b). By spanning a range of reasoning granularities, these tasks facilitate a systematic evaluation. A summarisation of each task and corpus is provided in Table 4.
Mathematical reasoning.
Mathematical expressions Valentino et al. (2023); Meadows et al. (2023) follow a well-defined syntactic structure and set of symbolic rules that are notoriously difficult for neural models. The dataset Meadows et al. (2023) includes seven human-annotated symbolic rules, encompassing operations such as differentiation and integration.
Syllogistic reasoning.
Syllogistic reasoning involves classical categorical logic, including four standard forms: Universal Affirmative (A) — “All A are B,” Universal Negative (E) — “No A are B,” Particular Affirmative (I) — “Some A are B,” and Particular Negative (O) — “Some A are not B.” The dataset encodes 24 valid syllogistic inference patterns, such as AA1 (Barbara).
Explanatory reasoning.
Explanatory sentences Jansen et al. (2018); Valentino et al. (2022) provide a semantically challenging yet sufficiently well-scoped scenario to evaluate the syntactic and semantic organisation of the space. Based on the EntailmentBank corpus Dalvi et al. (2021), Zhang et al. (2024b) defined ten inference types based on the transformation over Abstract Meaning Representation (AMR) graph Banarescu et al. (2013).
3.2 Rule-based Neural Tangent Kernel
Figure 2: Conceptual visualisation of rule separation in parametric space of encoder. We can quantify rule separation by computing the distance between gradient vectors associated with distinct rule categories.
This work aims to encode latent reasoning rules within a low-dimensional latent space to guide the LM decoder’s reasoning generation, leveraging a VAE framework. We frame the rule-based learning as learning the transformation from input premise(s) to output conclusion within the encoder. We formalise this framework by introducing a proposition grounded in the Neural Tangent Kernel (NTK) theory Jacot et al. (2018).
3.2.1 Latent Subspace Separation
Let $\mathcal{M}$ be a VAE model parameterised by $\theta=(\theta_{\text{enc}},\theta_{\text{dec}})$. Suppose the encoder can encode a set of symbolic inference rules $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. Then, under supervised training with rule annotations:
@@ 3.2.1 Latent Subspace Separation
This parametric separation directly leads to clearly delineated latent feature subspaces, each uniquely encoding a symbolic inference rule.
3.2.2 Connection to NTK Theory
NTK theory provides a rigorous theoretical framework for understanding neural network training dynamics by examining the kernel induced by gradients of the network’s outputs with respect to its parameters. It suggests that during gradient descent, the network effectively learns linear approximations within distinct functional subspaces where each approximating discrete symbolic reasoning rules in our task. Formally:
Step 1: formal setup.
Let $f_{\text{encode}}$ be the encoder function such that:
@@ 3.2.2 Connection to NTK Theory
where $x$ is the input premise(s), c is the conclusion, $\pi\in\Pi$ is the reasoning rule, and $\mathcal{Z}$ is the latent representation space of dimension $D$. The training objective is to minimise the expected loss:
$\mathcal{L}(\theta_{\text{enc}})=\mathbb{E}_{(\pi,x,c)\sim\mathcal{D}}\left[-\log p_{\theta}(\pi,x,c)\right]$ (3)
Step 2: NTK as rule subspaces.
Each rule $\pi$ is explicitly embedded as part of the model input. As a result, the model effectively learns a function $f_{\theta}(\pi,x,c)$. The function $f_{\theta}$ thus jointly depends on both the content of transformation from premise(s) to conclusion and the nature of the symbolic operation to be performed.
Within the NTK framework, the similarity between two input examples of the same inference type $\pi$ is captured by the NTK as follows:
@@
In here, $x$ represents $(x,c)$ pair for concision. $\nabla_{\theta}f_{\theta}(x,\pi)$ denotes the gradient of the model output with respect to its parameters, evaluated at the input $(x,\pi)$. This kernel quantifies how a parameter update from one input-output pair would affect another pair, conditioned on the shared rule.
According to NTK theory, the evolution of the model’s predictions under gradient descent training can be described by a linear kernel regression in the RKHS (Reproducing Kernel Hilbert Space) associated with $\Theta_{\pi}$. Specifically, the prediction at time $t$, $f_{t}(x,\pi)$, evolves as:
@@
Step 3: Disjointness via kernel independence.
For two different rules, $\pi_{i}\neq\pi_{j}$, we examine the relationship between their corresponding NTKs, $\Theta_{\pi_{i}}$ and $\Theta_{\pi_{j}}$. Specifically, we are interested in the interaction between the parameter gradients induced by inputs associated with different inference types.
Consider two data points $x$ and $x^{\prime}$, possibly corresponding to different premise pairs. The NTK entry for each rule is:
@@
When considering cross-rule similarities, we are interested in the inner product between the gradients for different rules:
@@
If the rules $\pi_{i}$ and $\pi_{j}$ encode fundamentally different reasoning operations (e.g., addition vs. subtraction in Math Derivation), the gradients with respect to $\theta$ for inputs labeled with $\pi_{i}$ and those labeled with $\pi_{j}$ will tend to point in different directions in parameter space of encoder.
@@
In the next section, we will introduce the practical VAE architecture and optimisation function.
@@ 4 Approach
Property 2: Semantic-level encoding. The latent space should also encode the lexical semantics of the $x$, accounting for fine-grained variations in reasoning patterns that arise from semantic differences in the inputs. For example, within the Math Derivation task, under differentiation, different inputs yield distinct transformations:
@@
[…]
By ensuring both properties, the latent space is capable of capturing both coarse-grained and fine-grained reasoning patterns.
@@
We adopt a Transformer-based VAE framework, employing two distinct encoders—both instantiated with BERT Devlin et al. (2018)—to model the prior and posterior Gaussian distributions. The posterior encoder is optimised to capture the reasoning rule transformation, thereby satisfying Property 1. Simultaneously, the prior encoder serves as a regulariser, encouraging the posterior to encode fine-grained sub-rule variations grounded in the lexical semantics of the input, thereby satisfying Property 2.
Concretely, given a target $(\pi,x)$ pair, we first retrieve a set of semantic similar examples $\{(\pi,x_{1},c_{1}),\dots,(\pi,x_{N},c_{N})\}$ using BM25 retrieval (N is 12 during the experiment). These examples are input to the posterior encoder, which learns a latent representation of the reasoning transformation by averaging the latent vectors of the retrieved instances. Geometrically, averaging the latent sample vectors positions the target $x$ at the centroid of the surrounding samples within the latent space, which is naturally aligned with how information is encoded in the latent space Zhang et al. (2024a, 2025).
The decoder then utilises this aggregated representation to generate the corresponding conclusion $c$ for the target triplet $(\pi,x,c)$. In parallel, the prior encoder processes the target $(\pi,x)$ pair to predict the rule $\pi$ within the latent space via a linear classifier. By minimising the Kullback–Leibler (KL) divergence between the posterior and prior distributions, the latent space is encouraged to satisfy both geometric properties.
@@
We adopt three strategies to inject the latent variable $z$ into the decoder (e.g., Qwen2.5 Qwen et al. (2025)). Each approach modifies the attention mechanism to incorporate information from the latent space.
@@
Following Zhang et al. (2024c), the latent vector $z$ is added to both the Key ($K$) and Value ($V$) matrices in the attention network:
$\text{softmax}\left(\frac{\text{Q}[\text{z}+\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z}+\text{V}]$ (11)
@@
Following Li et al. (2020), the latent vector $z$ is concatenated to the Key and Value matrices:
$\text{softmax}\left(\frac{\text{Q}[\text{z};\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z};\text{V}]$ (12)
@@
In this setup, the latent vector $z$ is added to the Query ($Q$) matrix:
$\text{softmax}\left(\frac{[\text{Q}+\text{z}]\text{K}^{\top}}{\sqrt{d}}\right)\text{V}$ (13)
@@
Finally, the model can be trained end-to-end via the evidence lower bound (ELBO) on the log-likelihood of the data $x$ Kingma and Welling (2013). To avoid the KL vanishing issue, we select the cyclical schedule to increase weights of KL $\beta$ from 0 to 1 Fu et al. (2019) and a KL thresholding scheme Li et al. (2019) that chooses the maximum between KL and threshold $\lambda$. The final objective function can be described as follows:
@@
[…]
@@
5 Empirical Analysis
5.1 Decoding Evaluation
@@
[…]
During inference, we omit $c$, allowing the model to generate the conclusion. To evaluate generation quality, we employ both the BLEU score Papineni et al. (2002) and accuracy (acc) as performance metrics. For the Math Reasoning task, in addition to the in-distribution test set, we also assess model performance on an out-of-distribution (OOD) test set, where the mathematical expressions are composed using a different set of variables.
@@
Table 2: Quantitative evaluation, where the base model is Qwen2.5-0.5B. Top two values are highlighed in blue and bold. We can observe that query_add leads to the best performance in general.
@@
Next, we evaluate the performance of VAE on downstream reasoning tasks, where the decoder is Qwen2.5-0.5B, the latent dimension is 32 following the same setup as Optimus Li et al. (2020). In addition, we include results from the base model trained via fine-tuning (denoted as FT) as well as inference using few-shot examples. The input format remains consistent with previous settings, with examples repeated accordingly. For few-shot selection, we employ BM25 to retrieve the most relevant samples.
As illustrated in Table 2, we can observe that injecting the latent space into the query can generally result in better performance compared with other setups. Intuitively, injecting reasoning information into the Query enables the model to more effectively retrieve the stored value Value from memory based on Key (Finding).
Furthermore, when the base model is trained using few-shot examples, its performance declines significantly. We observe that the model repeats to generate more examples. To ensure a fair comparison across methods, we did not apply any filtering to remove these redundant outputs (Finding).
Additionally, we evaluate the reasoning performance of the VAE model trained with and without BM25-based sample selection. In the absence of BM25, training samples are randomly drawn from the corpus. As presented in Table 3, integrating BM25-based selection during training yields enhanced reasoning performance, indicating the effectiveness of relevance-guided sampling. From a geometric perspective, BM25 retrieves samples that exhibit the highest lexical similarity, effectively selecting the nearest neighbors in the latent space (Finding).
In the following section, we evaluate the model’s capability to learn and represent latent reasoning rules within the latent space.

### v2

@@ Abstract
Incorporating explicit reasoning rules within the latent space of language models (LMs) offers a promising pathway to enhance generalisation, interpretability, and controllability. While current Transformer-based language models have shown strong performance on Natural Language Inference (NLI) tasks, they often rely on memorisation rather than explicit rule-based generalisation. This work investigates how human-interpretable reasoning rules can be explicitly encoded within LMs with the support of Language Variational Autoencoders (VAEs), as a mechanism for generative control. We propose a complete pipeline for learning reasoning rules within Transformer-based language VAEs. This pipeline encompasses three rule-based reasoning tasks, a supporting theoretical framework, and a practical end-to-end architecture. The experiment illustrates the following findings: Disentangled reasoning: Under explicit signal supervision, reasoning rules (viewed as functional mappings) can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space. Prior knowledge injection: injecting rule-based constraints into the Query enables the model to more effectively retrieve the stored Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into decoder-only language models. Moreover, we found that FFN layers are better than attention layers at preserving the separation of reasoning rules in the model’s parameters.
Introduction
@@
Encoding reasoning patterns as explicit rules within latent representations holds significant promise for enhancing the generalisation, interpretability, and controllability of neural models, with high downstream impact on AI safety and regulatory compliance (Bonnet and Macfarlane 2024; Yu, Chatzi, and Kissas 2025). Over recent years, Transformer-based language models (LMs) have achieved notable success across a variety of Natural Language Inference (NLI) tasks (Yang et al. 2024; Qwen et al. 2025). Nonetheless, a growing body of research has demonstrated that in many instances these models often rely on memorisation rather than generalisation and that rule-based control mechanisms fail to be fully enforced (Yan et al. 2025).
Therefore, this investigation focuses on the question: How can reasoning rules be explicitly encoded within the latent space of language models? In the context of this work, a reasoning rule is defined as an input–output pattern that reflects a specific transformation of inference behaviour. It is important to note that this study does not aim to explore rule composition or generalisation. Rather, the primary objective is to explicitly encode reasoning input–output constraints within the model’s latent space, targeting better interpretability and controllability in the latent space.
@@
Variational Autoencoders (VAEs) (Kingma and Welling 2013) provide a compelling framework for this direction, where the integration of prior distribution serves as an inductive bias, enabling the model to leverage existing knowledge and providing a principled way to incorporate domain constraints (Papamarkou et al. 2024). This work focuses on lightweight LMs (<1B<1B) as VAE decoders, allowing accessible evaluation of training dynamics, memory capacity, and model updates of the Transformer architecture (Zhong et al. 2025; Morris et al. 2025). Additionally, because our primary motivation is to study latent rule learning within the encoder latent space, we fully re-train the decoder. As a result, the geometrical properties of the latent space are not constrained by the model’s architecture, enabling a more controlled analysis of representation learning.
Accordingly, we propose a complete pipeline for learning NLI reasoning rules within Transformer-based language VAEs (Li et al. 2020; Zhang et al. 2024b):
@@
First, we focus on three distinct rule-based reasoning tasks, each characterised by unique syntax, inference patterns, and levels of granularity: Mathematical Reasoning (Meadows et al. 2023), Syllogistic Reasoning (Valentino et al. 2025), and Explanatory Reasoning (Zhang et al. 2024a).
@@
Second, following the Neural Tangent Kernel (NTK) theory (Jacot, Gabriel, and Hongler 2018), we claim that by supervising the inference rule information, inference rules can be encoded in the parametric space (i.e., weight matrix) of the encoder, which subsequently leads to the rule separation in the latent sentence space. This formalisation allows us to explicitly encode reasoning rules in the latent space.
@@
Third, we propose an end-to-end VAE architecture where the latent space can encode both coarse-grained reasoning rules and fine-grained reasoning patterns that arise from semantic differences in the inputs, and evaluate three ways to inject the latent space into the LM decoder by intervening in the attention network.
@@
We conduct extensive experiments to assess the effectiveness of rule encoding and reasoning generation capabilities which elicits the following results and findings:
@@
Disentangled Reasoning: Under explicit signal supervision, reasoning rules, as functional mappings, can be disentangled within the encoder’s parametric space. This separation results in distinct clustering of rules in the output feature space, suggesting the potential for employing rule-based NTK theory to better understand the training dynamics and internal geometry of gradient-based neural NLI model.
@@
Prior Knowledge Injection: Different latent space injection setups result in varying levels of reasoning performance. The optimal configuration is achieved by injecting the latent space into the Query of the attention network. Intuitively, injecting rule-based reasoning constraints into the Query enables the model to more effectively retrieve the stored Value from memory based on Key. This approach offers a simple method for integrating prior knowledge into LMs.
@@
Information Bottleneck: A case study on the mathematical reasoning task reveals a performance bottleneck in decoder-only LMs, where increasing the number of samples per operation fails to yield performance improvements beyond a certain threshold. Additionally, we observe that FFN networks are more effective than attention in maintaining the separation of reasoning rules within the learned parametric space. Both provide additional insights on understanding Transformer architectures through memorisation.
@@
To the best of our knowledge, this is the first study to explore the explicit encoding of NLI rules within language VAE latent spaces, targeting better latent space geometry to support rule-based inferences.
Related Work
@@
In this section, we review related work around two topics: rule-based representation learning and language VAEs, to highlight current research limitations and elucidate the motivation underlying our work.
@@
Encoder-only models have been employed for tasks such as mathematical operations, where the encoder learns structured transformations, as demonstrated by (Valentino et al. 2024). In addition, VAE-based approaches have shown promise in tasks requiring structured reasoning, such as program synthesis, where the goal is to generate programs that fulfil a specified task (Sun et al. 2018; Bonnet and Macfarlane 2024; van Krieken et al. 2025). Similarly, grammar-based approaches using VAEs have been applied to infer ordinary differential equation (ODE) formulas from data (Yu, Chatzi, and Kissas 2025). Moreover, Decoder-only models, particularly large language models (LLMs), leverage in-context learning by using demonstrations to infer and apply underlying reasoning patterns (Liu, Xing, and Zou 2023; Bhattamishra et al. 2024). Despite these advancements, relatively few studies have investigated rule-based learning in the context of NLI, where this study focuses on.
@@
Language VAEs have been widely applied in NLP tasks, such as style transfer tasks: modifying sentences with regard to markers of sentiment, formality, affirmation/negation (Shen et al. 2020; John et al. 2019; Bao et al. 2019a; Hu and Li 2021; Vasilakes et al. 2022; Gu et al. 2022; Liu et al. 2023; Gu et al. 2023), story generation (Fang et al. 2021), dialogue generation (Zhao, Zhao, and Eskenazi 2017), text paraphrasing (Bao et al. 2019b), and textual, syntactic, semantic representation learning domain, such as syntax disentanglement (Mercatali and Freitas 2021), semantic-syntax separation (Zhang et al. 2024b), semantic disentanglement (Carvalho et al. 2023; Zhang, Carvalho, and Freitas 2024), etc. Comparatively, in this work we focus on Natural Language Inference (NLI) with an emphasis on rule-based control.
@@
Rule-based Learning for NLI
Natural Language Inference Rules
We investigate three distinct types of reasoning tasks: mathematical derivation (Meadows et al. 2024), syllogistic reasoning (Valentino et al. 2025), and explanatory reasoning (Zhang et al. 2024a). Specifically, (1) Mathematical reasoning: Mathematical expressions (Valentino et al. 2023; Meadows et al. 2023) follow a well-defined syntactic structure and set of symbolic rules that are notoriously difficult for neural models. The dataset (Meadows et al. 2023) includes seven human-annotated symbolic rules, encompassing operations such as differentiation and integration. (2) Syllogistic reasoning: Syllogistic reasoning involves classical categorical logic, including four standard forms: Universal Affirmative (A) — “All A are B,” Universal Negative (E) — “No A are B,” Particular Affirmative (I) — “Some A are B,” and Particular Negative (O) — “Some A are not B.” The dataset encodes 24 valid syllogistic inference patterns, such as AA1 (Barbara). (3) Explanatory reasoning: Consists of material inferences with clearly defined explanatory sentence structures (Jansen et al. 2018; Valentino et al. 2022) providing a semantically challenging yet sufficiently well-scoped scenario to evaluate the syntactic and semantic organisation of the space. Based on the EntailmentBank corpus (Dalvi et al. 2021), Zhang et al. (2024a) we can define ten inference types based on these explanatory patterns, providing a diverse set of tasks which instantiate rule-based reasoning. A summarisation of each task and corpus is provided in Table 4 in the supplementary material.
Rule-based Neural Tangent Kernel
This work aims to encode latent reasoning rules within a low-dimensional latent space to guide the LM decoder’s reasoning generation, leveraging the VAE framework. We frame rule-based learning as learning the transformation from input premise(s) to output conclusion within the encoder. We formalise this framework by introducing a novel method grounded in the Neural Tangent Kernel (NTK) theory (Jacot, Gabriel, and Hongler 2018).
Latent subspace separation.
Let $\mathcal{M}$ be a VAE model parameterised by $\theta=(\theta_{\text{enc}},\theta_{\text{dec}})$. Suppose the encoder can represent a set of symbolic inference rules $\Pi=\{\pi_{1},\pi_{2},\dots,\pi_{n}\}$. Then, under supervised training with rule annotations:
@@
This parametric separation (Figure 2) directly leads to clearly delineated latent feature subspaces (Figure 3), each uniquely encoding a symbolic inference rule.
Connection to NTK theory.
NTK theory provides a rigorous theoretical framework for understanding neural network training dynamics by examining the kernel induced by gradients of the network’s outputs with respect to its parameters. It suggests that during gradient descent, the network effectively learns linear approximations within distinct functional subspaces, each approximating discrete symbolic reasoning rules within the target reasoning task. Formally, let $f_{\text{encode}}$ be the encoder function such that:
@@
where $x$ is the input premise(s), c is the conclusion, $\pi\in\Pi$ is the reasoning rule, and $\mathcal{Z}$ is the latent representation space of dimension $D$. Each rule $\pi$ is explicitly embedded as part of the model input. As a result, the model effectively learns a function $f_{\theta}(\pi,x,c)$. The function $f_{\theta}$ thus jointly depends on both the content of transformation from premise(s) to conclusion and the nature of the symbolic operation to be performed. Within the NTK framework, the similarity between two input examples of the same inference type $\pi$ is captured by the NTK as follows:
@@
where $x$ represents $(x,c)$ pair for concision. $\nabla_{\theta}f_{\theta}(x,\pi)$ denotes the gradient of the model output with respect to its parameters, evaluated at the input $(x,\pi)$. This kernel quantifies how a parameter update from one input-output pair would affect another pair, conditioned on the shared rule.
According to the NTK theory, the evolution of the model’s predictions under gradient descent training can be described by a linear kernel regression in the RKHS (Reproducing Kernel Hilbert Space) associated with $\Theta_{\pi}$. Specifically, the prediction at time $t$, $f_{t}(x,\pi)$, evolves as:
@@
For two different rules, $\pi_{i}\neq\pi_{j}$, we examine the relationship between their corresponding NTKs, $\Theta_{\pi_{i}}$ and $\Theta_{\pi_{j}}$. Specifically, we are interested in the interaction between the parameter gradients induced by inputs associated with different inference types. Given two data points $x$ and $x^{\prime}$, possibly corresponding to different premise pairs, the NTK entry for each rule is:
@@
When considering cross-rule similarities, we are interested in the inner product between the gradients for different rules.
@@
If the rules $\pi_{i}$ and $\pi_{j}$ encode fundamentally different reasoning operations (e.g., addition vs. subtraction in Math Derivation), the gradients with respect to $\theta$ for inputs labeled with $\pi_{i}$ and those labeled with $\pi_{j}$ will tend to point in different directions in parameter space of the encoder.
@@
In the next section, we will use this foundation to introduce the proposed VAE architecture and its supporting optimisation function.
@@
We adopt a Transformer-based VAE framework, employing two distinct encoders, both instantiated with BERT (Devlin et al. 2018), to model the prior and posterior Gaussian distributions. The posterior encoder is optimised to capture the reasoning rule transformation, thereby satisfying Property 1. Simultaneously, the prior encoder serves as a regulariser, encouraging the posterior to encode fine-grained sub-rule variations grounded in the lexical semantics of the input, thereby satisfying Property 2.
Concretely, given a target $(\pi,x)$ pair, we first retrieve a set of semantic similar examples $\{(\pi,x_{1},c_{1}),\dots,(\pi,x_{N},c_{N})\}$ using a retrieval function (e.g. BM25 and N=12). These examples are defined as inputs to the posterior encoder, which learns a latent representation of the reasoning transformation by averaging the latent vectors of the retrieved instances. Geometrically, averaging the latent sample vectors positions the target $x$ at the centroid of the surrounding samples within the latent space, which is naturally aligned with how information is encoded in the latent space (Zhang, Carvalho, and Freitas 2025). The decoder then uses this aggregated representation to generate the corresponding conclusion $c$ for the target triplet $(\pi,x,c)$. In parallel, the prior encoder processes the target $(\pi,x)$ pair to predict the rule $\pi$ within the latent space via a linear classifier. By minimising the Kullback–Leibler (KL) divergence between the posterior and prior distributions, the latent space is encouraged to satisfy both geometric properties.
@@
We adopt three strategies to inject the latent variable $z$ into the decoder (e.g., Qwen2.5 (Qwen et al. 2025)). Each approach modifies the attention mechanism to incorporate information from the latent space.
@@
Following (Zhang et al. 2024b), the latent vector $z$ is added to both the Key ($K$) and Value ($V$) matrices in the attention network: $\text{softmax}\left(\frac{\text{Q}[\text{z}+\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z}+\text{V}]$
@@
Following (Li et al. 2020), the latent vector $z$ is concatenated to the Key and Value matrices: $\text{softmax}\left(\frac{\text{Q}[\text{z};\text{K}]^{\top}}{\sqrt{d}}\right)[\text{z};\text{V}]$
@@
In this setup, the latent vector $z$ is added to the Query ($Q$) matrix: $\text{softmax}\left(\frac{[\text{Q}+\text{z}]\text{K}^{\top}}{\sqrt{d}}\right)\text{V}$
@@
Finally, the model can be trained end-to-end via the evidence lower bound (ELBO) on the log-likelihood of the data $x$ (Kingma and Welling 2013). To avoid the KL vanishing issue, we select the cyclical schedule to increase weights of KL $\beta$ from 0 to 1 (Fu et al. 2019) and a KL thresholding scheme (Li et al. 2019) that chooses the maximum between KL and threshold $\lambda$. The final objective function can be described as follows:
@@
$+\texttt{cls\_weight}\times\mathcal{L}_{\text{classifier}}$
@@
Empirical Analysis
Decoding Evaluation
@@
First, we evaluate the reasoning performance of different decoder-only models, including Qwen2 (Yang et al. 2024), Qwen2.5 (Qwen et al. 2025), GPT2 (Radford et al. 2019), and trimmed Llama3 (Grattafiori et al. 2024). Due to the scale of the dataset and limitations in computational resources, we restrict our fine-tuning to smaller models with fewer than 1 billion parameters. During training, the input format of decoder is described as: $\text{operation:}~\pi,~\text{premise:}~x,~\text{conclusion:}~c$. During inference, we omit $c$, allowing the model to generate the conclusion. All models are evaluated on the same testset. Further experimental details are provided in the supplementary material.
To evaluate generation quality, we employ both the BLEU score (Papineni et al. 2002) and accuracy (acc) as performance metrics. For the Math Reasoning task, in addition to the in-distribution test set, we also assess model performance on an out-of-distribution (OOD) test set, where the mathematical expressions are composed using a different set of variables.
@@
Table 2: Quantitative evaluation, where the base model is Qwen2.5-0.5B. Top two values are highlighed in bold and bold. Prior=False is the setup of VAE without trainable prior. We can observe that query_add leads to the best performance in general.
@@
Next, we evaluate the performance of the proposed VAE setting on downstream reasoning tasks, where the decoder is Qwen2.5-0.5B, the latent dimension is 32 following the same setup as Optimus (Li et al. 2020). In addition, we include results from the base model trained via fine-tuning (denoted as FT) as well as inference using few-shot examples. The input format remains consistent with previous settings, with examples repeated accordingly. For few-shot selection, we employ BM25 to retrieve the most relevant samples.
As illustrated in Table 2, we can observe that injecting the latent space into the query can generally result in better performance compared with base model and other setups. Intuitively, injecting reasoning information into the Query enables the model to more effectively retrieve the stored Value from memory based on the Key (Finding 1). In this setup, models with a trainable prior demonstrate improved performance on mathematical reasoning tasks, but not necessarily on other NLI tasks. This performance gain is attributed to the highly regular and syntactically consistent nature of the target mathematical expressions. In such contexts, retrieving structurally similar examples can effectively enhance model performance by reinforcing pattern recognition.
Furthermore, when the base model is trained using few-shot examples, its performance declines significantly. We observe that the model repeats to generate more examples. To ensure a fair comparison, we did not apply any filtering to remove these redundant outputs (Finding 2).
Additionally, we evaluate the performance of the VAE model trained with and without BM25-based sample selection. In the absence of the relevance function, training samples are randomly drawn from the corpus. In Table 3, integrating BM25-based selection during training yields enhanced performance, indicating the effectiveness of relevance-guided sampling. From a geometric perspective, BM25 retrieves samples that exhibit the highest lexical similarity, effectively selecting the nearest neighbours in the latent space (Finding 3).